RESONANCE-ENHANCED MULTIPHOTON IONIZATION SPECTROSCOPY OF CH_{3} and CD_{3}. TWO-PHOTON ABSORPTION SELECTION RULES AND ROTATIONAL LINE STRENGTHS OF THE ν_{3}− AND ν_{4}− ACTIVE VIBRONIC TRANSITIONS

K. CHEN, Department of Chemistry, National Sun Yat-sen University, Kaohsiung, Taiwan, ROC.

To explore the possibility of K -level resolved, 2+1 resonance-enhanced multiphoton ionization (REMPI) processes of the methyl radical, the two-photon absorption selection rules and rotational line strengths of the 3_{0}^{1} and 4_{0}^{1} vibronic bands of the transition np←X̃(n=3 or 4) were reported. Stringent selection rules, which were imposed upon these two-photon transitions, are the initial K′′ = 3p (p = 0, 1, 2, ...), ΔK = ±2, ΔU = ±3 and ΔN = 0, ±1, ±2 (O, P, Q, R, and S branches). The previously assigned 2_{2}^{2} vibronic band of the methyl radical should be studied by the REMPI with a better spectral resolution and analyzed by the newly derived two-photon absorption selection rules and rotational line strength formulas.